OpenStudy (pup02):

medal to anyone that helps Match each system of linear equations with the quadrant in which the solution lies. 3x + 2y = 14 x - 3y = -10 2x + 3y = -12 4x - 2y = -2 3x - 2y = 10 5x + 3y = 4 x + 2y = 8 4x + 6y =22

OpenStudy (pup02):

anyone there

OpenStudy (kayders1997):

Hi there!

OpenStudy (pup02):

can you help me

OpenStudy (kayders1997):

For these I would solve for x or y than do substitution for the variable you solved or, you could also do the elimination method, getting the x or y's to cancel and and than getting an answer for x or y and plugging that back in into one of those variables to find the other one

OpenStudy (kayders1997):

I can explain both ways for you on any number you want

OpenStudy (pup02):

can you help with all them plz

OpenStudy (kayders1997):

Sure, i'll do substitution for 2 and elimination for 2 if thats okay so you know two ways, elimination is my favorite personally

OpenStudy (pup02):

you will do all them in two ways

OpenStudy (kayders1997):

3x+2y=14 x-3y=-10 Would you like to figure out x or y? Me personally I would do x for the bottom because all you have to do is add 3y to both sides, but its your choice

OpenStudy (pup02):

I just want to get the question thx tho

OpenStudy (kayders1997):

Thats how you have to do it though, otherwise it will take way to long to try and find the coordinates that work for both equations

OpenStudy (kayders1997):

Unless you get lucky

OpenStudy (pup02):

ok thx this is all new to me

OpenStudy (kayders1997):

Have you learned the substitution or elimination method?

OpenStudy (kayders1997):

You would have to cancel one of the variables for elimination its much easier than substitution and takes way less time

OpenStudy (pup02):

I think but when you solve the problems I understand

OpenStudy (kayders1997):

Okay, if you wanna try number one if you think you can let me know what you get and I'll tell you what I get if you don't know where to start i'll help

OpenStudy (pup02):

I don't know how to do this its why I put it in this site

OpenStudy (kayders1997):

Okay so 3x+2y=14 X-3y=-10 So you would have to mutiple the bottom by -3 the whole function, because than the x's would cancel so you would have 2y=14 on the top on and -3(-3y)=-3(-10) So than you will have 2y=14 on the top and 9y=30 and you add those values so you will have 11y=44 than divide by 11 so y=4 and than plug 4 in for 4 so 3x+2(4)=14 so than 3x+8=14 subtract 8 3x=6 x=2 so your coordinate is (2, 4) so the coordinate is in the first quadrant

OpenStudy (mathmale):

Making use of Kayder's suggestion, and multiplying the 2nd equation by -1, we get first equation: 3x + 2y = 14 -3x+ 9y = 30 ------------- Please add these two equations together, column by column. What's left? ( ? ) y = ( ? )

OpenStudy (pup02):

thx are you guy and girl going to help with the others? plz

OpenStudy (mathmale):

I'd be glad to help after you've started each problem and gone as far as you can before getting stuck. At that point, ask questions!

OpenStudy (kayders1997):

Same, I wanna be a math teacher so I would love to help you

OpenStudy (amr979):

i want to help you pup02

OpenStudy (pup02):

ok you all can help I just don't get this sorry

OpenStudy (kayders1997):

If it was on paper would that help more?

OpenStudy (pup02):

idk

OpenStudy (amr979):

iam sorry , iam not very good in english so, please tell me what do you want exactly for the previous equations

OpenStudy (kayders1997):

You camt add pictures shoot

OpenStudy (pup02):

what are the other three

OpenStudy (kayders1997):

Do you still need help on the first part?

OpenStudy (kayders1997):

Where you have to multiply

OpenStudy (kayders1997):

To eliminate a variable

OpenStudy (pup02):

I got the first one I just need you to do the other three problems

OpenStudy (kayders1997):

2x+3y=-12 4x-2y=-2 what would you have to multiply to the top to get opposites for x's

OpenStudy (kayders1997):

Opposites signs, but same number so basically get -4x

OpenStudy (amr979):

you want to know what are x and y equal at all the three equations?

OpenStudy (pup02):

what quadrant the solution is in

OpenStudy (pup02):

so it would be in the third quadrant

OpenStudy (amr979):

yes i understand

OpenStudy (kayders1997):

That depends on what you get for x and y if you get positive for both you get quadrant 1 if x is negitive and y positive than its quad 2 if they are both negitive than quad 3 if y is negitive and x is positive than y is negitive than it is quad 4

OpenStudy (kayders1997):

But you first have to do that elimination method

OpenStudy (amr979):

so pup let me solve it for you . to know which quadrant we must get the values of x and y so in the first problem x=3.2 and y=4.4 so it will be at quadrant 1

OpenStudy (pup02):

that is for the first quastion

OpenStudy (amr979):

yes

OpenStudy (kayders1997):

No

OpenStudy (kayders1997):

Plug it back i to both equtions you dont get those answers

OpenStudy (amr979):

for the equations 3x+y=14 x-3y=-10

OpenStudy (pup02):

ok what are the other three

OpenStudy (kayders1997):

The answer is x=2 y=4

OpenStudy (pup02):

ok what is the answer for the second answer.

OpenStudy (amr979):

yes you are right because i have seen the first equation 3x+y=14

OpenStudy (amr979):

so the answers are for the first problem x=2 y=4 an the quadrant 1 for the second problem x=-15/8 y=-11/4 and the quardant is 3

OpenStudy (pup02):

ok that's really is great thanks for helping me

OpenStudy (pup02):

what is the third question

OpenStudy (amr979):

if you want any help contact me anytime

OpenStudy (amr979):

the third is x=2 y=-2 in the quadrant 4

OpenStudy (pup02):

thank you

OpenStudy (kayders1997):

2x + 3y = -12 4x - 2y = -2 Okay for this equation you would multiply the top by -2 so than -2(2x+3y equal -12) so than you will have -4x+-6y=24 and the bottom equation would stay the same so now you add -4x+4x=0 And -6y +-2y =-8y and 24+-2=22 so than -8y=22 so than it will be 22/-8 which can be reduced to -11/4 and than plug that into y and than it will be 2x=+3(-11/4) =-12 so 2x-33/4=-12 so add 33/4 so 2x=2x=-15/4 than divide by 2 so x=-15/8 so both x and y are negitive so than you will be in quadrant 4

OpenStudy (kayders1997):

quadrant 3 my bad

OpenStudy (pup02):

haha thank you all